# Quantum State Preparation for Probability Distributions with Mirror Symmetry Using Matrix Product States

### v2

@@ Abstract
Quantum circuits for loading probability distributions into quantum states are essential subroutines in quantum algorithms used in physics, finance engineering, and machine learning. The ability to implement these with high accuracy in shallow quantum circuits is a critical issue. We propose a novel quantum state preparation method for probability distribution with mirror symmetry using matrix product states. By considering mirror symmetry, our method reduces the entanglement of probability distributions and improves the accuracy of approximations by matrix product states. As a result, we improved the accuracy by two orders of magnitude over existing methods using matrix product states. Our approach, characterized by a shallow quantum circuit primarily comprising nearest-neighbor qubit gates and linear scalability with qubit count, is highly advantageous for noisy quantum devices. Also, our experimental findings reveal that the approximation accuracy in tensor networks depends heavily on the bond dimension, with minimal reliance on the number of qubits. Our method is experimentally demonstrated for a normal distribution encoded into 10 and 20 qubits on a real quantum processor.
1 Introduction
@@
It is known that the preparation of quantum states for arbitrary functions without error requires an exponential depth in quantum circuits[9, 10]. In recent years, several methods have been proposed for state preparation using shallower quantum circuits by tolerating a degree of error. In quantum machine learning, a proposed method involves creating quantum circuits to generate target probability distributions, utilizing quantum Generative Adversarial Networks with parameterized quantum circuits as Generators[11, 12]. Methods utilizing approximations via Fourier transforms have also been proposed[13, 14]. Furthermore, approaches utilizing tensor networks have also been actively researched in recent times[15, 16, 17, 18].
We propose a new state preparation method utilizing the mirror symmetry of probability distributions and tensor networks[16, 17]. We note that functions with less entanglement can be approximated more accurately when using matrix product states. Specifically, in the normal distribution, the portion to the left of the mean monotonically increases and possesses a structure with less entanglement. Therefore, using matrix product states, we could convert the left half of the normal distribution into quantum circuits with higher accuracy. Subsequently, we add a quantum circuit that “duplicates” by the Hadamard gate and inverts this half by CNOT gates. This quantum circuit has enabled us to prepare the entire normal distribution from the quantum circuit that prepares the left half of the normal distribution. Our method achieves a two-order-of-magnitude improvement in normal distribution accuracy compared to methods using existing matrix product states. In addition, we have also confirmed that accuracy improvements can be similarly achieved with distributions with mirror symmetry other than the normal distribution. Our method features a quantum circuit in which most parts consist of gates between nearest-neighbor qubits. The circuit exhibits an extremely shallow circuit depth that scales linearly with the number of qubits. If the topology of the quantum computer being used is linear, the fact that CNOT gates act only on nearest-neighbor qubits implies that SWAP gates are not necessary. This makes it a beneficial property for noisy quantum devices. We experimentally showed that when using tensor networks, the approximation accuracy heavily depends on the bond dimension of the tensor network and is mainly independent of the number of qubits. Finally, we demonstrated that we lord the normal distribution into 10 and 20 qubits on a real quantum processor using our designed quantum circuits, achieving high fidelity.
2 Tensor Networks to Quantum Circuits
@@
where $G_{jklm}$ represents the element of the gate $G^{[i]}$, and $j$ represents the state of an ancilla qubit. Recalling that $A^{[i]}$ is an isometry, we can readily extend $G^{[i]}$ to be a unitary gate. Then, we determine the elements of the gate $G^{[1]}$ from the first tensor $A^{[1]}$ as follows:
@@
where $G_{jklm}$ represents the element of the gate $G^{[1]}$, and $j$ and $k$ represents the state of ancilla qubits. Similarly, we can extend $G^{[1]}$ as a unitary gate. Using these gates $G^{[i]}$, the $U_{{\rm MPD}}$ can be constructed, as shown in Figure 1(a), as follows:
@@
where $G^{[i]}$ ($1\leq i<n$) acts on the $i$-th and $i$+1-th qubits, and $G^{[n]}$ acts on the $n$-th qubit.
Figure 2: The Kullback-Leibler divergence between the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=0.01)$ (or the left half of normal distribution) and the probability distributions represented by each MPS (10 sites) with limited bond dimension.
Next, we describe the method when the bond dimension is limited to $\chi>2$. The following approach is based on reducing the bond dimension by using $U_{{\rm MPD}}$. For instance, we consider the state |$\ket{\Psi^{(4)}}$ when $\chi=4$. We approximate $\chi=2$, and construct $U_{{\rm MPD}}^{(2)}$ using the aforementioned method. Then, using this $U_{{\rm MPD}}^{(2)}$, we disentangle |$\ket{\Psi^{(4)}}$. In other words, we consider $U_{{\rm MPD}}^{(2)}\ket{\Psi^{(4)}}$ to be MPS with approximately $\chi=2$ and approximate $U_{{\rm MPD}}^{(2)}\ket{\Psi^{(4)}}$ again to $\chi=2$ to construct $U_{{\rm MPD}}^{(4)}$. Therefore, it can be considered that $U_{{\rm MPD}}$ of the state |$\ket{\Psi^{(4)}}$ can be approximated as follows:
@@
3 Quantum Circuit for Normal Distribution Utilizing Symmetry
In this section, we show that we can prepare quantum circuits with higher accuracy by considering the mirror symmetry of normal distributions.
3.1 Entanglement and Symmetry
{quantikz}
@@ 3.1 Entanglement and Symmetry
[…]
Figure 3: Quantum circuit considering the mirror symmetry of the normal distribution (five qubits). (a) If a quantum circuit $U_{{\rm MPD}}$ exists that creates the left half of a normal distribution, it can be extended to the right half. (b) At the point of (1), a quantum state with the left half of the normal distribution as its probability amplitude is created by the quantum circuit $U_{{\rm MPD}}$. (c) At the point of (2), the left half of the normal distribution is duplicated by applying the Hadamard gate to an ancilla qubit. (d) At the point of (3), by applying CNOT gates with the ancilla qubit as the control qubit to other qubits, the left half of the right-side normal distribution is transformed into the right half.
With a lower bond dimension, as shown in Figure 2, it is generally difficult to reproduce any arbitrary probability distribution with high accuracy using MPS. Lower bond dimensions cause larger errors because the bond dimension is related to the entanglement of the quantum state. This implies that probability distributions with less entanglement will likely have smaller errors than those with larger entanglement.
We consider that segmenting the probability distribution to reduce entanglement could allow us to prepare the distribution with greater precision. For example, while the entanglement measure of the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=0.01)$ is $0.237\times 10^{-2}$, the entanglement measure of the left half of the normal distribution is $8.31\times 10^{-6}$. Therefore, the left half of the normal distribution can be approximated with greater precision than the normal distribution. As shown in Figure 2, it was found that the Kullback-Leibler(KL) divergence is more than two orders of magnitude smaller at lower bond dimensions. Remark that for the entanglement measure, we used the following measure of multiple qubit entanglement proposed in previous research [22]:
@@
$\imath_{j}(b)\ket{b_{1}\ldots b_{n}}=\delta_{bb_{j}}\ket{b_{1}\ldots\widehat{b_{j}}\ldots b_{n}}.$ (14)
We have discovered that we can reproduce the normal distribution with high precision if we only consider half of it, but we need to prepare the remaining right half as well. Here, the mirror symmetry of the normal distribution becomes significant. We propose a quantum circuit that can “copy” the right half of the normal distribution from the left half, as shown in Figure 3. Therefore, we can create a high-precision normal distribution from its high-accuracy left half.
3.2 Result
@@
Finally, we demonstrated our methods with a quantum processor ibm_torino using 10 and 20 qubits. We converted and learned the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=1)(\mathrm{min}:-4\sqrt{3},\mathrm{max}:4\sqrt{3})$ and measured the final state with 100,000 shots for 10-qubits experiment and 3,000,000 shots for 20-qubits experiment. Several error mitigation methods were utilized by the sampler in qiskit runtime primitives. M3 readout error mitigation [25] was performed for the 10-qubit experiment but not for the 20-qubit one because it timed out. Dynamical decoupling [26, 27] was also enabled.
@@ 3.3 Other Functions
Figure 6: Our methodology applied to (a) the Lorentzian function(min: -5, max: 5), and (b) the Student’s t-distribution($\nu$ = 2, min: -10, max: 10).
4 Conclusion
We proposed a method that utilizes mirror symmetry to achieve higher precision in preparing the normal distribution when converting from tensor networks to quantum circuits. The depth of this quantum circuit scales linearly with the number of qubits, even when considering the maximum bond dimension. Our method improved the accuracy (KL divergence) by two orders of magnitude compared to existing methods using tensor networks. We showed how bond dimension and the number of qubits influence the precision when preparing the normal distribution using tensor networks. Specifically, we showed that increasing the bond dimension improves accuracy, and the number of qubits has a negligible impact on accuracy. Furthermore, we confirmed that our method improves accuracy even in distributions with mirror symmetry, in addition to the normal distribution. Finally, we implemented our quantum circuit that loads the normal distribution on a real quantum processor. We conducted experiments with 10 qubits and 20 qubits quantum circuits, demonstrating that we could achieve high fidelity in each instance.
4.1 Comparison with Prior Works
@@
[…]
@@
When compared to existing methods using tensor networks[16, 17], the significant difference with our approach lies in the accuracy and the CNOT gates between (2) and (3) in Figure 3(a). Regarding accuracy, it is as described above. The CNOT gate acts only between nearest-neighbor qubits in existing tensor network methods. This characteristic in quantum computers with a linear topology means that SWAP gates are not required. This characteristic, which saves SWAP gates, benefits noisy quantum devices without error correction capabilities. In our method, as shown before point (1) in Figure 3(a), the CNOT gate is applied only between nearest-neighbor qubits. However, the CNOT gates between points (2) and (3) in Figure 3(a) act between the ancilla qubit and all qubits, necessitating SWAP gates. In this respect, our method is at a disadvantage compared to existing methods.
@@
The Grover–Rudolph (GR) state preparation [9] can prepare exact quantum states, but the number of gates and the circuit depths required increases exponentially with the number of qubits. The Kitaev–Webb method [28] allows for similarly accurate state preparation with polynomial circuit depth. However, it has been pointed out that for fewer than 15 qubits, it requires a greater circuit depth than the GR method [29]. Additionally, in the Fourier Series Loader (FSL) method [14] using an m-qubit Fourier transform, the GR state preparation is employed to lord the Fourier coefficients, necessitating a circuit depth of $2^{m}$. To prepare quantum states accurately using the FSL method, a moderately large size for m is necessary (for example, around $m=6$, or more). Consequently, this necessitates a significantly large CNOT circuit depth in practice. Moreover, the inverse quantum Fourier transform, which utilizes approximately $\mathcal{O}(n^{2})$ two-qubit gates, also requires a considerably sizeable quantum circuit. A significant disadvantage of these methods is the requirement to use CNOT gates between all qubits, which necessitates many SWAP gates. Since implementing a logical SWAP gate requires three CNOT gates, these methods necessitate a more significant number of CNOT gates for implementation. This requirement presents a challenge for executing noisy quantum devices requiring logical SWAP gates. In contrast, while our method is less accurate, it excels over these methods in terms of shallow CNOT circuit depth and fewer SWAP gate requirements.
@@
Finally, we compare our approach with methods that use machine learning [11, 12, 18]. These methods’ advantage is their ability to determine circuit depth and gate count based on the number of parameters. If accurate approximations can be achieved with fewer parameters, these methods can implement shallower quantum circuits compared to our method and other existing methods that depend on the number of qubits. However, finding a suitable arrangement of parameters and the parameters themselves is challenging in practice. These studies have been limited to discovering quantum circuits with an accuracy on the order of $\mathcal{O}(10^{-4})$. The fact that the CNOT gates in the quantum circuit act only between nearest-neighbor qubits is beneficial.
@@
In summary, our method offers better precision than approaches using tensor networks or machine learning and maintains a quantum circuit depth comparable to these methods. While it is less precise than GR state preparation methods, our method significantly benefits from a much shallower quantum circuit. Therefore, our method is compelling for noisy quantum devices with lower CNOT gate fidelity[30, 31].
[…]
Circuit Gate Classical Methods Depths Counts Accuracy Optimization Our method ($\chi=2$) $\mathcal{O}(n)$ $\mathcal{O}(n)$ $\mathcal{O}(10^{-5})$$\times$ Our method ($\chi>2$) $\mathcal{O}(\log{\chi}+n)$ $\mathcal{O}(\log{\chi}+n)$ $\mathcal{O}(10^{-7})$$\times$ Tensor Network ($\chi=2$)[16, 17] $\mathcal{O}(n)$ $\mathcal{O}(n)$ $\mathcal{O}(10^{-3})$$\times$ Tensor Network ($\chi>2$)[16, 17] $\mathcal{O}(\log{\chi}+n)$ $\mathcal{O}(\log{\chi}+n)$ $\mathcal{O}(10^{-5})$$\times$ GR State Preparation[9] $\mathcal{O}(2^{n})$ $\mathcal{O}(2^{n})$ exact $\times$ Kitaev-Webb[28] $\mathcal{O}(poly(n))$ $\mathcal{O}(poly(n))$ any $\times$ FSL[14] $\mathcal{O}(n^{2},2^{m})$ $\mathcal{O}(n^{2},2^{m})$ any $\times$ Quantum GAN[11, 12] $\mathcal{O}(p)$ $\mathcal{O}(np)$ $\mathcal{O}(10^{-3})$bigcirc TN + Machine Learning[18] $\mathcal{O}(p)$ $\mathcal{O}(np)$ $\mathcal{O}(10^{-4})$bigcirc
4.2 Future Work

### v3

# Quantum State Preparation for Probability Distributions with Reflection Symmetry Using Matrix Product States

@@ Abstract
Quantum circuits for loading probability distributions into quantum states are essential subroutines in quantum algorithms used in physics, finance engineering, and machine learning. The ability to implement these with high accuracy in low-depth quantum circuits is a critical issue. We propose a novel quantum state preparation method for probability distribution with reflection symmetry using matrix product states. By considering reflection symmetry, our method reduces the entanglement of probability distributions and improves the accuracy of approximations by matrix product states. As a result, we improved the accuracy by two orders of magnitude over existing methods using matrix product states. Our approach, characterized by linear scalability with qubit count, is highly advantageous for noisy quantum devices. Also, our demonstration results reveal that the approximation accuracy in tensor networks depends heavily on the bond dimension, with minimal reliance on the number of qubits. Our method is demonstrated for a normal distribution encoded into 10 and 20 qubits on a real quantum processor.
I Introduction
@@
It is known that the preparation of quantum states for arbitrary functions without error requires an exponential depth in quantum circuits[9, 10]. In recent years, several methods have been proposed for state preparation using lower depth quantum circuits by tolerating a degree of error. In quantum machine learning, a proposed method involves creating quantum circuits to generate target probability distributions, utilizing quantum Generative Adversarial Networks with parameterized quantum circuits as Generators[11, 12]. Methods utilizing approximations via Fourier transforms have also been proposed[13, 14]. Furthermore, approaches utilizing tensor networks have also been actively researched in recent times[15, 16, 17, 18].
We propose a new state preparation method utilizing the reflection symmetry of probability distributions and tensor networks[16, 17]. We note that functions with less entanglement can be approximated more accurately when using matrix product states. Specifically, in the normal distribution, the portion to the left of the mean monotonically increases and possesses a structure with less entanglement. Therefore, using matrix product states, we could convert the left half of the normal distribution into quantum circuits with higher accuracy. Subsequently, we add a quantum circuit that “duplicates” by the Hadamard gate and inverts this half by CNOT gates. This quantum circuit has enabled us to prepare the entire normal distribution from the quantum circuit that prepares the left half of the normal distribution. Our method achieves a two-order-of-magnitude improvement in normal distribution accuracy compared to methods using existing matrix product states. In addition, we have also confirmed that accuracy improvements can be similarly achieved with distributions with reflection symmetry other than the normal distribution. Our method features a quantum circuit in which most parts consist of gates between nearest-neighbor qubits. The circuit exhibits an extremely low circuit depth that scales linearly with the number of qubits. This makes it a beneficial property for noisy quantum devices. We showed that when using tensor networks, the approximation accuracy heavily depends on the bond dimension of the tensor network and is mainly independent of the number of qubits. Finally, we demonstrated that we lord the normal distribution into 10 and 20 qubits on a real quantum processor using our designed quantum circuits, achieving high fidelity.
II Tensor Networks to Quantum Circuits
@@
where $G_{jklm}$ represents the matrix element of the gate $G^{[i]}$, with $j$ denoting the computational basis state of the reflection qubit, $k$ corresponding to the input qubit from the previous gate, and $l$ and $m$ representing the computational basis states of the output qubits. Recalling that $A^{[i]}$ is an isometry, we can readily extend $G^{[i]}$ to be a unitary gate. Then, we determine the elements of the gate $G^{[1]}$ from the first tensor $A^{[1]}$ as follows:
@@
where $G_{jklm}$ represents the element of the gate $G^{[1]}$, and $j$ and $k$ represents the computational basis state of the reflection qubit. Similarly, we can extend $G^{[1]}$ as a unitary gate. Using these gates $G^{[i]}$, the $U_{{\rm MPD}}$ can be constructed, as shown in Figure 1(a), as follows:
@@
where $G^{[i]}$ ($1\leq i<n$) acts on the $i$-th and ii+1-th qubits, and $G^{[n]}$ acts on the $n$-th qubit.
Figure 2: The Kullback-Leibler divergence between the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=0.01)$ $(\mathrm{min}:-0.5,\mathrm{max}:0.5)$ (or the left half of normal distribution) and the probability distributions represented by each MPS (10 sites) with limited bond dimension.
[…]
@@
III Quantum Circuit for Normal Distribution Utilizing Symmetry
In this section, we show that we can prepare quantum circuits with higher accuracy by considering the reflection symmetry of normal distributions.
III.1 Entanglement and Symmetry
@@
Figure 3: Quantum circuit considering the reflection symmetry of the normal distribution (five qubits). (a) If a quantum circuit $U_{{\rm MPD}}$ exists that creates the left half of a normal distribution, it can be extended to the right half. Let us call the highest qubit that is used for this reflection the reflected qubit. (b) At the point of (1), a quantum state with the left half of the normal distribution as its probability amplitude is created by the quantum circuit $U_{{\rm MPD}}$. (c) At the point of (2), the left half of the normal distribution is duplicated by applying the Hadamard gate to a reflection qubit. (d) At the point of (3), by applying CNOT gates with the reflection qubit as the control qubit to other qubits, the left half of the right-side normal distribution is transformed into the right half.
In general, it is difficult to accurately reproduce an arbitrary probability distribution using an MPS with a low bond dimension, and we confirmed this in the case of the normal distribution in Figure 2. Lower bond dimensions cause larger errors because the bond dimension is related to the entanglement of the quantum state. This implies that probability distributions with less entanglement will likely have smaller errors than those with larger entanglement.
We consider that segmenting the probability distribution to reduce entanglement could allow us to prepare the distribution with greater precision. For example, while the entanglement measure of the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=0.01)$ is $0.237\times 10^{-2}$, the entanglement measure of the left half of the normal distribution is $8.31\times 10^{-6}$. Therefore, it is expected that the half-normal distribution can be approximated with low entanglement, and indeed, as shown in Figure 2, we found that the approximation accuracy for the half-normal distribution saturates at a lower bond dimension. Remark that for the entanglement measure, we used the following measure of multiple qubit entanglement proposed in previous research [22]:
@@
[…]
We have discovered that we can reproduce the normal distribution with high precision if we only consider half of it, but we need to prepare the remaining right half as well. Here, the reflection symmetry of the normal distribution becomes significant. We propose a quantum circuit that can “copy” the right half of the normal distribution from the left half, as shown in Figure 3. Therefore, we can create a high-precision normal distribution from its high-accuracy left half.
The quantum circuit in Figure 3 can, in a sense, be regarded as a form of $U_{MPD}$ that successfully extracts the entanglement structure of a probability distribution with reflection symmetry. This can be seen as an example of incorporating prior knowledge to extract the entanglement structure of a quantum state—something that could not be achieved by constructing $U_{MPD}$ solely through MPS and the Disentangler.
III.2 Result
@@
Finally, we demonstrated our methods with a quantum processor ibm_torino using 10 and 20 qubits. We converted and learned the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=1)(\mathrm{min}:-4\sqrt{3},\mathrm{max}:4\sqrt{3})$ and measured the final state with 100,000 shots for 10-qubit demonstration and 3,000,000 shots for 20-qubits demonstration. Several error mitigation methods were utilized by the sampler in qiskit runtime primitives. M3 readout error mitigation [25] was performed for the 10-qubit demonstration but not for the 20-qubit one because it timed out. Dynamical decoupling [26, 27] was also enabled.
@@ III.3 Other Functions
Figure 6: Our methodology applied to (a) the Lorentzian function(min: -5, max: 5), and (b) the Student’s t-distribution($\nu$ = 2, min: -10, max: 10).The max bond dimension refers to $\chi=2^{n-1}$.
IV Conclusion
We proposed a method that utilizes reflection symmetry to achieve higher precision in preparing the normal distribution when converting from tensor networks to quantum circuits. The depth of this quantum circuit scales linearly with the number of qubits, even when considering the maximum bond dimension. Our method improved the accuracy (KL divergence) by two orders of magnitude compared to existing methods using tensor networks. We showed how bond dimension and the number of qubits influence the precision when preparing the normal distribution using tensor networks. Specifically, we showed that increasing the bond dimension improves accuracy, and the number of qubits has a negligible impact on accuracy. Furthermore, we confirmed that our method improves accuracy even in distributions with reflection symmetry, in addition to the normal distribution. Finally, we implemented our quantum circuit that loads the normal distribution on a real quantum processor. We performed demonstrations using 10- and 20-qubit quantum circuits, achieving high fidelity in each case.
[…]
Circuit Gate Classical
Methods Depths Counts Accuracy Optimization
Our method ($\chi=2$) $\mathcal{O}(n)$ $\mathcal{O}(n)$ $\mathcal{O}(10^{-5})$ $\times$
Our method ($\chi>2$) $\mathcal{O}(\log{\chi}+n)$ $\mathcal{O}(\log{\chi}+n)$ $\mathcal{O}(10^{-7})$ $\times$
Tensor Network ($\chi=2$)[16, 17] $\mathcal{O}(n)$ $\mathcal{O}(n)$ $\mathcal{O}(10^{-3})$ $\times$
Tensor Network ($\chi>2$)[16, 17] $\mathcal{O}(\log{\chi}+n)$ $\mathcal{O}(\log{\chi}+n)$ $\mathcal{O}(10^{-5})$ $\times$
GR State Preparation[9] $\mathcal{O}(2^{n})$ $\mathcal{O}(2^{n})$ exact $\times$
Kitaev-Webb[28] $\mathcal{O}(poly(n))$ $\mathcal{O}(poly(n))$ any $\times$
FSL[14] $\mathcal{O}(n^{2},2^{m})$ $\mathcal{O}(n^{2},2^{m})$ any $\times$
[…]
IV.1 Comparison with Prior Works
@@
We provide a rough estimate of the circuit depth and gate count for our method. It is known that the $G^{[i]}$ gates, created using tensor networks, can be constructed from two CNOT gates and several single-qubit gates [17]. Furthermore, we demonstrated that considering a bond dimension of $\chi=2048$, which corresponds to $L=11$, allows us to achieve the accuracy on $\mathcal{O}(10^{-7})$. Then, the quantum circuit $U_{{\rm MPD}}$ has CNOT gate depth of 2(n2)2(n-2) when the bond dimension is $\chi=2$, and has CNOT gate depth of 2(n2)+(L1)2{(n-2)+(L-1)} when the bond dimension is $\chi=2^{L}$. For instance, if $n=20$ and $L=11$, the CNOT circuit depth would be 60. Our method requires not only the $U_{{\rm MPD}}$ but also CNOT gates from a reflection qubit, as is shown between points (2) and (3) in Figure 3(a). Therefore, the total CNOT circuit depth for the entire setup becomes either 2(n2)+n12(n-2)+n-1 when $\chi=2$, and 2(n2)+L+n12{(n-2)+L}+n-1 when $\chi=2^{L}$.
@@
When compared to existing methods using tensor networks[16, 17], the significant difference with our approach lies in the accuracy and the CNOT gates between (2) and (3) in Figure 3(a). Regarding accuracy, it is as described above. The CNOT gate acts only between nearest-neighbor qubits in existing tensor network methods. This characteristic in quantum computers with a linear topology means that SWAP gates are not required. This characteristic, which saves SWAP gates, benefits noisy quantum devices without error correction capabilities. In our method, as shown before point (1) in Figure 3(a), the CNOT gate is applied only between nearest-neighbor qubits. However, the CNOT gates between points (2) and (3) in Figure 3(a) act between the reflection qubit and all qubits, necessitating SWAP gates. In this respect, our method is at a disadvantage compared to existing methods.
@@
The Grover–Rudolph (GR) state preparation [9] can prepare exact quantum states, but the number of gates and the circuit depths required increases exponentially with the number of qubits. The Kitaev–Webb method [28] allows for similarly accurate state preparation with polynomial circuit depth. However, it has been pointed out that for fewer than 15 qubits, it requires a greater circuit depth than the GR method [29]. Additionally, in the Fourier Series Loader (FSL) method [14] using an m-qubit Fourier transform, the GR state preparation is employed to lord the Fourier coefficients, necessitating a circuit depth of $2^{m}$. To prepare quantum states accurately using the FSL method, a moderately large size for m is necessary (for example, around $m=6$, or more). Consequently, this necessitates a significantly large CNOT circuit depth in practice. Moreover, the inverse quantum Fourier transform, which utilizes approximately $\mathcal{O}(n^{2})$ two-qubit gates, also requires a considerably sizeable quantum circuit. A significant disadvantage of these methods is the requirement to use CNOT gates between all qubits, which necessitates many SWAP gates. Since implementing a logical SWAP gate requires three CNOT gates, these methods necessitate a more significant number of CNOT gates for implementation. This requirement presents a challenge for executing noisy quantum devices requiring logical SWAP gates. In contrast, while our method is less accurate, it excels over these methods in terms of low CNOT circuit depth and fewer SWAP gate requirements.
@@
Finally, we compare our approach with methods that use machine learning [11, 12, 18]. These methods’ advantage is their ability to determine circuit depth and gate count based on the number of parameters. If accurate approximations can be achieved with fewer parameters, these methods can implement lower depth quantum circuits compared to our method and other existing methods that depend on the number of qubits. However, finding a suitable arrangement of parameters and the parameters themselves is challenging in practice. These studies have been limited to discovering quantum circuits with an accuracy on the order of $\mathcal{O}(10^{-4})$. The fact that the CNOT gates in the quantum circuit act only between nearest-neighbor qubits is beneficial.
@@
In summary, our method offers better precision than approaches using tensor networks or machine learning and maintains a quantum circuit depth comparable to these methods. While it is less precise than GR state preparation methods, our method significantly benefits from a much lower depth quantum circuit. Therefore, our method is compelling for noisy quantum devices with lower CNOT gate fidelity [30, 31].
IV.2 Application
Our method efficiently prepares reflection-symmetric probability distributions—such as the normal distribution, Student’s $t$-distribution, and Lorentzian function—which frequently serve as initial states in quantum algorithms across finance and physics.
IV.2.1 Finance
In financial models, the normal distribution is often used to describe asset returns, portfolio risk, and market uncertainty, particularly in models based on Brownian motion. The t-distribution is also commonly employed due to its heavier tails compared to the normal distribution. Quantum algorithms for option pricing, portfolio optimization, and risk estimation—such as those based on amplitude estimation or Monte Carlo methods—frequently require the initialization of quantum states reflecting such distributions [32, 33, 34].
IV.2.2 Simulation of Physical Systems
It is known that, when performing physical simulations of quantum field theory on a quantum computer, one uses a normal distribution as the initial state [35, 36]. In fluid-dynamics simulations using the Quantum Lattice Boltzmann Method, a normal distribution is also used as the initial state [37, 38].
IV.2.3 Subroutine for State Preparation
Our method can also be employed as a subroutine for state preparation. Specifically, techniques have been proposed that use the Lorentz function to approximate various target distributions [39]; among these, one can replace the Lorentz function preparation with the Lorentz function generated by our method.
IV.3 Future Work